# ReactEMG: Zero-Shot, Real-time Intent Detection from sEMG via Masked Modeling

### v5

@@ Abstract
Surface electromyography (sEMG) signals show promise for effective human-machine interfaces, particularly in rehabilitation and prosthetics. However, challenges remain in developing systems that respond quickly and reliably to user intent, across different subjects and without requiring time-consuming calibration. In this work, we propose a framework for EMG-based intent detection that addresses these challenges. We cast intent detection as per-timestep segmentation of continuous sEMG streams, assigning labels as gestures unfold in real time. We introduce a masked modeling training strategy that aligns muscle activations with their corresponding user intents, enabling rapid onset detection and stable tracking of ongoing gestures. In evaluations against baseline methods, using metrics that capture both accuracy and stability for device control, our approach achieves state-of-the-art performance in zero-shot conditions. These results demonstrate its potential for wearable robotics and next-generation prosthetic systems. Our project website, video, code, and dataset are available at: https://reactemg.github.io/
@@ 1 Introduction
Learning-based methods are increasingly central to human-machine interfaces (HMI), enabling users to control a range of robotic systems through interactive modalities such as wearable sensors. Among these modalities, surface electromyography (sEMG) stands out for its non-invasive ability to detect muscle activity at the skin’s surface. Direct sEMG measurement of neuromuscular signals has supported non-clinical applications—including control of multi-fingered grippers [1, 2] and drones [3]—as well as clinical uses such as diagnosis of neuromuscular conditions and prosthesis control [4, 5].
@@
Despite these advances, interpreting EMG for device control remains challenging when models are run online to operate a device continuously in real-time. In this paper, we use the term real-time to refer to an online setting with two key properties: (i) low latency, where the delay from muscle activation to device action is small enough to feel immediate; and (ii) stability, where intent predictions remain consistent during sustained ”hold” periods instead of flickering between classes. This real‑time setting differs from offline gesture recognition, which classifies an entire movement after its completion and can thus exploit full temporal observability from gesture onset to offset. In contrast, real‑time control must infer intent from partial observations as gestures unfold, handle successive transitions with variable dwell times, and produce per‑timestep outputs suitable for continuous device operation. These distinctions are illustrated in Figure 1.
@@
Figure 1: Design space for real‑time EMG intent detection. Approaches that rely on full segment observability (top row) wait for an offset signal or a window long enough to cover the entire gesture before making a prediction. They typically assume each gesture is bounded by ”Relax” or ”Rest” states, or they segment the region that contains the non-rest gesture beforehand. Since dwell times vary, these methods introduce uncontrolled latency and does not match continuous use. Even dense variants (top‑right) inherit this dependence on known boundaries. With partial observability and a single label (left column), supervision becomes inconsistent when a window straddles a gesture transition—different assignment rules label nearly identical inputs differently—delaying switches and causing flicker that then requires smoothing. Our formulation (bottom‑right) treats intent detection as streaming segmentation: from the observed window it produces dense per‑timestep predictions, enabling immediate prediction and unambiguous supervision at transitions.
@@
More specifically, we use latency to refer to the delay from a change in muscle activity to the first correct change in the model’s output that would actuate the device. In rehabilitation, low latency matters because building continuous, strong associations between human exertion and observable movement is key for achieving therapeutic outcomes [6, 7, 8] and low latency in closing the human motion loop is needed for therapeutic benefits [9, 10]. More broadly, HMI applications require low latency because operators have limited patience for input-response delays—wait too long and they will attempt different control inputs to compensate [11, 12, 13], leading to unstable robot behaviors [14, 15].
@@
Once an intent is established, stability becomes equally important. We use this term to refer to prediction constancy during holds, with no flicker or spurious class switches. Consider a hand prosthesis assisting a user to grasp a mug: after closing, the system must maintain a ”closed” prediction throughout the hold. Even brief flicker to ”open” can cause accidental release, with practical and safety consequences.
@@
Real‑time EMG intent detection therefore requires both rapid, reliable transition detection and robust maintenance of the active intent. The most commonly used metric for such systems—per-time-step accuracy—aggregates performance over an entire window, which can mask critical effects pertaining to these requirements. Therefore, in addition to per‑timestep accuracy, and to better capture latency and stability, we introduce transition accuracy—a single metric that reflects how quickly the model reacts at transitions and how consistently it maintains the active intent across the hold. Unlike conventional offline metrics, transition accuracy is designed to more accurately and faithfully capture aspects of performance that are most relevant to online device control behaviors.
@@
Beyond real‑time performance, another major challenge is zero-shot generalization: a model that works across new users immediately, without subject‑specific calibration or fine‑tuning. This ”plug‑and‑play” capability is appealing because collecting per‑user EMG data via repetitive gestures is time‑consuming and physically demanding, especially for people with limited mobility. Achieving zero‑shot performance is challenging, however, as EMG varies significantly across individuals and sessions due to anatomical differences, physiological state, and day‑to‑day drift [16].
@@
To achieve real‑time, zero‑shot performance, we present ReactEMG, a combination of a novel learning model, training framework, dataset, and evaluation metric for EMG intent detection. We formulate intent detection as a segmentation problem, in which we predict an action label (i.e., user intent) at each timestep within a continuous EMG stream. For training, we provide both EMG signals and corresponding actions as input, and adopt a masked‑modeling strategy that selectively masks portions of these sequences. By requiring the model to reconstruct the missing segments, we leverage local supervision that anchors muscle activations to user intent, enabling robust signal-action alignment even under imprecise ground‑truth labels. Evaluated with both per-time-step and transition accuracy metrics, our approach achieves state‑of‑the‑art zero‑shot performance on unseen subjects, without any subject‑specific calibration. In summary, our contributions are as follows:
@@
  • State-of-the-art performance for real-time control. We target intent prediction for device control that is responsive (fast switching and tolerant of rapid successive changes), stable (flicker-free holds), and strictly online (no hindsight segmentation or offline assumptions). Using transition accuracy, a stringent metric that requires a correct switch within a short window and maintenance of the new intent with zero mislabels, ReactEMG outperforms baselines on EMG-EPN-612 and on our newly collected dataset. Our model also matches or exceeds leading sequence- and convolution-based alternatives on standard per-timestep accuracy and latency metrics.
@@ 2 Related Work
Gesture recognition via sEMG has long been investigated as a control signal for HMI, particularly for prosthetic devices, assistive robotics, and rehabilitation systems [17, 18, 19, 20, 21, 22, 23, 24]. Earlier approaches predominantly relied on extracting handcrafted features from the sEMG signals followed by shallow classifiers such as LDAs [25] or SVMs [26, 27, 28]. Recent research has shifted toward deep learning paradigms that better exploit sEMG’s spatiotemporal structure, including convolutional neural networks (CNNs) for spatial filtering across electrode arrays [29, 30], recurrent networks (RNNs/LSTMs) for modeling temporal dynamics [31, 32], and transformer architectures to capture long‑range, cross‑channel interactions via self‑attention [33, 34].
@@
While strong results are commonly reported on popular datasets (e.g., NinaPro [35] and CAPG‑MYO [36]), they do not transfer directly to real‑time control. Most evaluations operate on pre‑segmented windows (from labeled onset to offset) or on gesture‑only (non‑relax) excerpts, then aggregate decisions per gesture via majority voting or smoothing—assuming full observability of the entire gesture. This assumption conflicts with continuous control, where observability is partial and gesture end times are unknown. Moreover, methods that predict a single label per window introduce label ambiguity when a window straddles a gesture transition, delaying gesture switches and masking prediction flicker. Figure 1 illustrates some of these differences.
@@
Several works aim for real‑time use, make assumptions and use evaluation protocol that present hurdles for real-time integration. Montazerin et al. [34] report fast recognition with small windows and low latency, but the input is pre‑segmented to exclude the ”rest” state, unrealistic for online use. Furthermore, the ”fused” variant of their model relies on motor‑unit decomposition performed in hindsight, which the paper acknowledges prevents real‑time deployment. Azhiri et al. [31] report high performance on a window-level RNN classifiers, but do not quantify streaming properties such as responsiveness of successive transitions or hold‑time flicker, and exclude the ”rest” state, limiting online applicability. De Silva et al. [37] report high accuracy with fast inference, but predictions are made only at detected onsets and held between onsets, with a 2-second post-prediction refractory period that blocks any subsequent transitions. The pipeline also assumes that flexion is followed by its corresponding extension, which limits applicability to more arbitrary action sequences.
@@
Reducing the per‑user calibration burden has also been a long‑standing goal in sEMG gesture recognition. Most ”subject‑agnostic” approaches still require few‑shot adaptation—via self‑supervision [38], meta‑learning [39], generative augmentation [40], or loss design and normalization strategies [41]. These methods improve cross‑user robustness but remain calibration‑dependent and are typically validated on segmented windows. Zero‑shot generalization has gained momentum with large, diverse open-source datasets. Eddy et al. [42] show that training on the EMG‑EPN‑612 dataset improves cross‑user performance under zero-shot settings. However, their protocol (by design) targets discrete gesture recognition that assumes full gesture observability.
@@
In parallel with academia, industry has begun to explore zero‑shot generalization at scale. Meta has released two resources that advance cross-user modeling: emg2qwerty (108 users; 346 hours) for typing and emg2pose (193 users; 370 hours) for hand pose estimation [43, 44]. Both demonstrate promising zero‑shot results on held‑out users in offline evaluation. While emg2pose describes an internal buffered streaming variant of their model, the end‑to‑end setup depends on a proprietary wristband and the model is not open-source, precluding assessment of real‑time latency and stability.
@@ 3.1 Architecture
Although segmentation captures continuous changes in muscle activation, the EMG signal itself remains non-stationary and highly variable across time. Even within a single recording session, different gestures can induce distinct activation patterns, and these patterns may fluctuate due to electrode shift, fatigue, or user variability. Inspired by recent advances in multimodal LLMs and ”any-to-any” frameworks [45, 46, 47], we hypothesize that explicitly conditioning EMG representations on the corresponding action label—allowing the model to cross-attend between the two—can mitigate these challenges. To this extent, we propose an encoder-only transformer for EMG and action modeling. Unlike conventional methods that treat EMG signals and their corresponding action labels as an input-output pair, our model treats EMG and intent as two distinct input modalities. We construct a multimodal sequence that includes EMG data followed by the corresponding intent data. The model’s objective is to reconstruct the original unmasked sequence—both EMG and intent—at all timesteps. Figure 2 provides an overview of the proposed architecture.
@@
Next, we apply span masking to the projected EMG and intent embeddings. Following a protocol similar to RoBERTa’s [48] dynamic masking, we randomly mask out contiguous timesteps in either or both modalities and regenerate these masks at each training epoch, thereby exposing different regions of the same sequence to the model over time. After masking, we add modality-specific encoding vectors and positional encodings to both the EMG and action tokens, ensuring each token can be distinguished not only by its position in time but also by the modality it belongs to. These positional encodings are shared across both modalities to enforce consistent temporal alignment.
@@ 3.3 Real-Time Inference
Real-time inference does not come for free. Unlike offline evaluation, an online system must decide from a live stream under tight latency and compute budgets, with only partial observability and no oracle boundaries. Consequently, design choices and pipeline assumptions that are harmless offline quickly become untenable and unrealistic at run time: requiring known onset/offset boundaries (full-segment observability) [42, 34], removing the rest/relax state [34], relying on preprocessing statistics computed offline [31], producing a single label per window even when it straddles a transition [49, 33], or imposing action‑sequence rules [37].
@@
Our approach introduces two sources of latency. First, the label at time $t$ is issued after ingesting data up to $t+\ell$, introducing a delay of up to $\ell$ in responding to changes in intent. This delay is smaller than typical human reaction time to visual or verbal stimuli [50, 51, 52]. Second, the hold period prevents the model from reacting to intent changes faster than 10 Hz. We consider it highly unlikely that a human subject would modulate hand pose or intent at such a rate.
@@ 3.4 Datasets
Capturing the diversity of EMG signals is essential for robust zero-shot generalization. To this end, we aggregate four open‑source sEMG datasets [53, 54, 55, 56] recorded with a single Myo armband at 200 Hz and whose gestures can be mapped to a common six‑gesture set: hand relax, hand open, hand close, wave-in, wave-out, and pinch. We do not include other datasets so that sampling rate, sensor configuration, and gesture taxonomy remain compatible with our setting.

### v6

# ReactEMG: Stable, Low-Latency Intent Detection from sEMG via Masked Modeling

@@ Abstract
Surface electromyography (sEMG) signals show promise for effective human-machine interfaces, particularly in rehabilitation and prosthetics. However, challenges remain in developing systems that respond quickly to user intent, produce stable flicker-free output suitable for device control, and work across different subjects without time-consuming calibration. In this work, we propose a framework for EMG-based intent detection that addresses these challenges. We cast intent detection as per-timestep segmentation of continuous sEMG streams, assigning labels as gestures unfold in real time. We introduce a masked modeling training strategy that aligns muscle activations with their corresponding user intents, enabling rapid onset detection and stable tracking of ongoing gestures. In evaluations against baseline methods, using metrics that capture accuracy, latency and stability for device control, our approach achieves state-of-the-art performance in zero-shot conditions. These results demonstrate its potential for wearable robotics and next-generation prosthetic systems. Our project website, video, code, and dataset are available at: https://reactemg.github.io/
@@ 1 Introduction
Learning-based methods are increasingly central to human-machine interfaces (HMI), enabling users to control a range of robotic systems through interactive modalities such as wearable sensors. Among these modalities, surface electromyography (sEMG) stands out for its non-invasive ability to detect muscle activity at the skin’s surface. Direct sEMG measurement of neuromuscular signals has supported non-clinical applications—including control of multi-fingered grippers [Loskutova2024, Meeker2022] and drones [Doshi2021]—as well as clinical uses such as diagnosis of neuromuscular conditions and prosthesis control [Tchimino2024, reciprocal2025].
@@
Despite these advances, interpreting EMG for device control remains challenging when models are run online to operate a device continuously in real-time. In this paper, we use the term “real-time” to refer to an online setting with two key properties: (i) low latency, where the delay from muscle activation to device action is small enough to feel immediate; and (ii) stability, where intent predictions remain consistent during sustained “hold” periods instead of flickering between classes. This real-time setting differs from offline gesture recognition, which classifies an entire movement after its completion and can thus exploit full temporal observability from gesture onset to offset. In contrast, real-time control must infer intent from partial observations as gestures unfold, handle successive transitions with variable dwell times, and produce per‑timestep outputs suitable for continuous device operation. Prior works have largely emphasized offline evaluation, with relatively few studies reporting online deployment; even when online performance is reported, latency and stability are rarely quantified, and functional task designs often obscure model-specific effects and prevent systematic comparisons across different models. These distinctions are illustrated in Figure 1.
@@
Figure 1: Design space for real-time EMG intent detection. Approaches that rely on full segment observability (top row) wait for an offset signal or a window long enough to cover the entire gesture before making a prediction. They typically assume each gesture is bounded by “Relax” or “Rest” states, or they segment the region that contains the non-rest gesture beforehand. Since dwell times vary, these methods introduce uncontrolled latency and does not match continuous use. Even dense variants (top‑right) inherit this dependence on known boundaries. With partial observability and a single label (left column), supervision becomes inconsistent when a window straddles a gesture transition—different assignment rules label nearly identical inputs differently—delaying switches and causing flicker that then requires smoothing. Our formulation (bottom‑right) treats intent detection as streaming segmentation: from the observed window it produces dense per‑timestep predictions, enabling immediate prediction and unambiguous supervision at transitions.
@@
More specifically, we use latency to refer to the delay from a change in muscle activity to the first correct change in the model’s output that would actuate the device. In rehabilitation, low latency matters because building continuous, strong associations between human exertion and observable movement is key for achieving therapeutic outcomes [Leech2022motorlearning, Micera2020neurotech, Hogan2006engagement] and low latency in closing the human motion loop is needed for therapeutic benefits [Cisnal2021robhand, Schwartz2016closeloopmovement]. More broadly, HMI applications require low latency because operators have limited patience for input-response delays—wait too long and they will attempt different control inputs to compensate [Beech2025jitter, Scholcover2018behaviorlatency, Du2024delaycompensation], leading to unstable robot behaviors [Neumeier2019teleoperation, Khasawneh2019latency].
@@
Once an intent is established, stability becomes equally important. We use this term to refer to prediction constancy during holds, with no flicker or spurious class switches. Consider a hand prosthesis assisting a user to grasp a mug: after closing, the system must maintain a “closed” prediction throughout the hold. Even brief flicker to “open” can cause accidental release, with practical and safety consequences.
@@
Real-time EMG intent detection therefore requires both rapid, reliable transition detection and robust maintenance of the active intent. The most commonly used metric for such systems—per-time-step accuracy—aggregates performance over an entire window, which can mask critical effects pertaining to these requirements. Therefore, in addition to per‑timestep accuracy, we also focus on response latency and, in order to better capture both latency and stability, we also introduce transition accuracy—a single metric that reflects how quickly the model reacts at transitions and how consistently it maintains the active intent across the hold. Unlike conventional offline metrics, transition accuracy is designed to more accurately and faithfully capture aspects of performance that are most relevant to online device control behaviors.
@@
Beyond real-time performance, another major challenge is zero-shot generalization: a model that works across new users immediately, without subject‑specific calibration or fine‑tuning. This “plug‑and‑play” capability is appealing because collecting per‑user EMG data via repetitive gestures is time‑consuming and physically demanding, especially for people with limited mobility. Achieving zero‑shot performance is challenging, however, as EMG varies significantly across individuals and sessions due to anatomical differences, physiological state, and day‑to‑day drift [DeLuca1997].
@@
To achieve real-time, stable zero‑shot performance, we present ReactEMG, a combination of a novel learning model, training framework, dataset, and evaluation metric for EMG intent detection. We formulate intent detection as a segmentation problem, in which we predict an action label (i.e., user intent) at each timestep within a continuous EMG stream. For training, we provide both EMG signals and corresponding actions as input, and adopt a masked‑modeling strategy that selectively masks portions of these sequences. By requiring the model to reconstruct the missing segments, we leverage local supervision that anchors muscle activations to user intent, enabling robust signal-action alignment even under imprecise ground‑truth labels. Evaluated with both per-time-step and transition accuracy metrics, our approach achieves state‑of‑the‑art zero‑shot performance on unseen subjects, without any subject‑specific calibration. In summary, our contributions are as follows:
@@
  • State-of-the-art performance for low-latency control. We target intent prediction for device control that is responsive (fast switching and tolerant of rapid successive changes), stable (flicker-free holds), and strictly online (no hindsight segmentation or offline assumptions). Using transition accuracy, a stringent metric that requires a correct switch within a short window and maintenance of the new intent with zero mislabels, ReactEMG outperforms baselines on EMG-EPN-612 and on our newly collected dataset. Our model also matches or exceeds leading sequence- and convolution-based alternatives on standard per-timestep accuracy and latency metrics.
@@ 2 Related Work
Gesture recognition via sEMG has long been investigated as a control signal for HMI, particularly for prosthetic devices, assistive robotics, and rehabilitation systems [CHANG1996529, Khokhar2010, Gordleeva2020, Xiong2024, Hargrove2013, Cimolato2022, Ahkami2023, Leonardis2015]. Earlier approaches predominantly relied on extracting handcrafted features from the sEMG signals followed by shallow classifiers such as LDAs [Antuvan2019] or SVMs [crawford2005real, Tepe2022, JaramilloYnez2020]. Recent research has shifted toward deep learning paradigms that better exploit sEMG’s spatiotemporal structure, including convolutional neural networks (CNNs) for spatial filtering across electrode arrays [BaronaLpez2024, Betthauser2020], recurrent networks (RNNs/LSTMs) for modeling temporal dynamics [Azhiri2021, Simo2019], and transformer architectures to capture long‑range, cross‑channel interactions via self‑attention [Zabihi2023, Montazerin2023].
@@
While strong results are commonly reported on popular datasets (e.g., NinaPro [Pizzolato2017] and CAPG‑MYO [Dai2021]), they do not transfer directly to real‑time control. Most evaluations operate on pre‑segmented windows (from labeled onset to offset) or on gesture‑only (non‑relax) excerpts, then aggregate decisions per gesture via majority voting or smoothing—assuming full observability of the entire gesture. This assumption conflicts with continuous control, where observability is partial and gesture end times are unknown. Moreover, methods that predict a single label per window introduce label ambiguity when a window straddles a gesture transition, delaying gesture switches and masking prediction flicker. Figure 1 illustrates some of these differences.
@@
Several works aim for real‑time use, make assumptions and use evaluation protocol that present hurdles for real-time integration. Montazerin et al. [Montazerin2023] report fast recognition with small windows and low latency, but the input is pre‑segmented to exclude the ”rest” state, unrealistic for online use. Furthermore, the ”fused” variant of their model relies on motor‑unit decomposition performed in hindsight, which the paper acknowledges prevents real‑time deployment. Azhiri et al. [Azhiri2021] report high performance on a window-level RNN classifiers, but do not quantify streaming properties such as responsiveness of successive transitions or hold‑time flicker, and exclude the ”rest” state, limiting online applicability. De Silva et al. [Silva2020] report high accuracy with fast inference, but predictions are made only at detected onsets and held between onsets, with a 2-second post-prediction refractory period that blocks any subsequent transitions. The pipeline also assumes that flexion is followed by its corresponding extension, which limits applicability to more arbitrary action sequences.
@@
Reducing the per‑user calibration burden has also been a long‑standing goal in sEMG gesture recognition. Most ”subject‑agnostic” approaches still require few‑shot adaptation—via self‑supervision [Duan2023], meta‑learning [LaRotta2024], generative augmentation [Xu2025], or loss design and normalization strategies [Xue2023]. These methods improve cross‑user robustness but remain calibration‑dependent and are typically validated on segmented windows. Zero‑shot generalization has gained momentum with large, diverse open-source datasets. Eddy et al. [Eddy2024] show that training on the EMG‑EPN‑612 dataset improves cross‑user performance under zero-shot settings. However, their protocol (by design) targets discrete gesture recognition that assumes full gesture observability.
@@
In parallel with academia, industry has begun to explore zero‑shot generalization at scale. Meta has released two resources that advance cross-user modeling: emg2qwerty (108 users; 346 hours) for typing and emg2pose (193 users; 370 hours) for hand pose estimation [emg2qwerty_neurips2024, salteremg2pose]. Both demonstrate promising zero‑shot results on held‑out users in offline evaluation. While emg2pose describes an internal buffered streaming variant of their model, the end‑to‑end setup depends on a proprietary wristband and the model is not open-source, precluding assessment of real‑time latency and stability.
@@ 3.1 Architecture
Although segmentation captures continuous changes in muscle activation, the EMG signal itself remains non-stationary and highly variable across time. Even within a single recording session, different gestures can induce distinct activation patterns, and these patterns may fluctuate due to electrode shift, fatigue, or user variability. Inspired by recent advances in multimodal LLMs and ”any-to-any” frameworks [tang2023anytoany, zhan-etal-2024-anygpt, nextgpt], we hypothesize that explicitly conditioning EMG representations on the corresponding action label—allowing the model to cross-attend between the two—can mitigate these challenges. To this extent, we propose an encoder-only transformer for EMG and action modeling. Unlike conventional methods that treat EMG signals and their corresponding action labels as an input-output pair, our model treats EMG and intent as two distinct input modalities. We construct a multimodal sequence that includes EMG data followed by the corresponding intent data. The model’s objective is to reconstruct the original unmasked sequence—both EMG and intent—at all timesteps. Figure 2 provides an overview of the proposed architecture.
@@
Next, we apply span masking to the projected EMG and intent embeddings. Following a protocol similar to RoBERTa’s [zhuang-etal-2021-robustly] dynamic masking, we randomly mask out contiguous timesteps in either or both modalities and regenerate these masks at each training epoch, thereby exposing different regions of the same sequence to the model over time. After masking, we add modality-specific encoding vectors and positional encodings to both the EMG and action tokens, ensuring each token can be distinguished not only by its position in time but also by the modality it belongs to. These positional encodings are shared across both modalities to enforce consistent temporal alignment.
@@ 3.3 Real-Time Inference
Real-time inference does not come for free. Unlike offline evaluation, an online system must decide from a live stream under tight latency and compute budgets, with only partial observability and no oracle boundaries. Consequently, design choices and pipeline assumptions that are harmless offline quickly become untenable and unrealistic at run time: requiring known onset/offset boundaries (full-segment observability) [Eddy2024, Montazerin2023], removing the rest/relax state [Montazerin2023], relying on preprocessing statistics computed offline [Azhiri2021], producing a single label per window even when it straddles a transition [Lee2021, Zabihi2023], or imposing action‑sequence rules [Silva2020].
@@
Our approach introduces two sources of latency. First, the label at time $t$ is issued after ingesting data up to $t+\ell$, introducing a delay of up to $\ell$ in responding to changes in intent. This delay is smaller than typical human reaction time to visual or verbal stimuli [Woods2015, Jain2015, Arafat2022Comparison]. Second, the hold period prevents the model from reacting to intent changes faster than 10 Hz. We consider it highly unlikely that a human subject would modulate hand pose or intent at such a rate.
@@ 3.4 Datasets
Capturing the diversity of EMG signals is essential for robust zero-shot generalization. To this end, we aggregate four open‑source sEMG datasets [epn_dataset, roshambo_dataset, Kanoga2021, mangalore_dataset] recorded with a single Myo armband at 200 Hz and whose gestures can be mapped to a common six‑gesture set: hand relax, hand open, hand close, wave-in, wave-out, and pinch. We do not include other datasets so that sampling rate, sensor configuration, and gesture taxonomy remain compatible with our setting.